OpenStudy (anonymous):

Solve for x: 2x^2 - 4x - 14 = 0 Explain the steps??

OpenStudy (anonymous):

2x^2 = 2*2 = 4x

OpenStudy (anonymous):

I'm using the quadratic formula; now I'm at x=4+/- sqrt 4240/4, what do I do now

OpenStudy (anonymous):

240****

OpenStudy (anonymous):

4x-4x-14=0

OpenStudy (anonymous):

tell me is this is right then... 2x2-4x-14=0 4x-4x-14=0 but then no solution because it's 0x-14=0???

OpenStudy (anonymous):

0x-14=0 so then add 14 0x = 14 So x would be undefined

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

Did that help?

OpenStudy (phi):

\[2x^2-4x-14=0\] a= 2, b= -4, c= -14 use the quadratic formula \[ x= \frac{1}{2a} \left( -b ± \sqrt{b^2 - 4 ac}\right) \] \[ x= \frac{1}{2\cdot 2} \left( -(-4) ± \sqrt{(-4)^2 - 4 \cdot 2 \cdot -14}\right) \\x= \frac{1}{4}\left( 4 ± \sqrt{16 +112}\right) \\ x= \frac{1}{4}\left( 4 ± \sqrt{128}\right) \] can you finish ?